OpenStudy (anonymous):

Jill, Grace, and Billy shared pennies in the ratio 2 : 4 : 4. They had a total of 120 pennies in all. How many pennies did Jill receive? 2 pennies 24 pennies 48 pennies 60 pennies

OpenStudy (anonymous):

r u done yet?

OpenStudy (texaschic101):

2:4:4 Jill -- 20% of 120 = .20 * 120 = 24 Grace -- 40% of 120 = .40 * 120 = 48 Billy -- 40% of 120 = 40 * 120 = 48 check... 24 + 48 + 48 = 120 120 = 120 (correct) Jill got 24

OpenStudy (anonymous):

The ratio of the number of coins collected by Billy to the number of coins collected by Grace is 6 : 9. When they combined their coins, they had a total of 300 coins. How many coins did Billy collect? 50 coins 60 coins 120 coins 180 coins

OpenStudy (anonymous):

next one

OpenStudy (texaschic101):

not sure about this one...sorry

OpenStudy (anonymous):

ill post it publicly

OpenStudy (texaschic101):

ok...sorry